# On the Self-awareness of Large Reasoning Models’ Capability Boundaries

### v2

@@ Appendix B System Prompt of BoostAbstention
Appendix C Confident and Uncertain Expressions
Figure 7 shows the confident and uncertain expression that frequently appear in the reasoning process. The expressions are summarized manually and checked by LLM. In the experiment, when matching these expressions, the Uncertain Expression is matched first, followed by the Confident Expression. Additionally, both the position of the first token of an expression and the token length of the entire expression are considered.
@@ Appendix C Confident and Uncertain Expressions
Appendix D Capability Boundary Revealed by Hidden States
@@
Figure 8 visualizes the classification of solvable and unsolvable questions of all LRMs, along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train–test splits. These results demonstrate that the features of capability boundaries are strongly encoded in the hidden states of the last input token, and is valid across different LRMs.
Appendix E Unproductive Reasoning
@@
Figure 9 and Figure 10, demonstrate representative unproductive reasoning phenomena in large language models, including repetitive looping and error accumulation, which impair their capacity for sound and consistent reasoning in complex problem-solving scenarios. They respectively show how models get stuck in repetitive deductions (Figure 9) and how errors compound during formula application and result derivation (Figure 10). In mathematical problem reasoning, the former presents repeated deduction of the same uncertain conclusion, while the latter sees continuous error accumulation due to flawed initial methods.

### v3

@@ Appendix D System Prompt of BoostAbstention
Appendix E Additional Analysis of Capability Boundary Signals in Hidden States
In Section 5, we show that eventually solved and unsolved questions can be separated using signals extracted from the hidden states of the last input token. However, these signals may partially correlate with question difficulty rather than reflecting a purely intrinsic notion of capability boundary.
In this section, we do not aim to establish a clean separation between capability boundary signals and question difficulty. Instead, we provide additional analyses showing that the hidden state signals identified in our setup remain informative under stricter per-benchmark and cross-benchmark settings.
An intuition from human exams. One possible intuition comes from human test taking. In an exam with limited time and many difficult problems, skilled test-takers may quickly identify questions that are unlikely to be solved within the available budget and avoid wasting excessive time on them. This judgment may come either from recognizing that a question resembles problems they have repeatedly struggled with in the past (analogous to our Monitor strategy), or from taking a few initial steps and realizing that they lack confidence to proceed (analogous to our Monitor strategy). We present this analogy only as intuition, not as evidence of human-like metacognitive awareness in LRMs. Motivated by this intuition, we further conduct per-benchmark and cross-benchmark analyses as follows.
| Dataset | Model | LDA | LR |
| HLE | GPT-oss-20B | 89.3 | 90.3 |
| R1-Distill-Qwen3-8B | 94.9 | 96.4 | |
| R1-Distill-Qwen3-32B | 96.4 | 96.4 | |
| QwQ-32B | 93.4 | 96.4 | |
| GSM8k | GPT-oss-20B | 80.3 | 83.3 |
| R1-Distill-Qwen3-8B | 70.1 | 71.6 | |
| R1-Distill-Qwen3-32B | 85.2 | 89.0 | |
| QwQ-32B | 87.9 | 87.1 |
Per-benchmark test. We provide in Table 6 the accuracy of separating unsolvable questions from solvable ones using LDA and LR classifiers trained and tested on splits from the same dataset. Specifically, for HLE and GSM8K, we randomly split each dataset into an 80% training set and a 20% test set (with balanced solvable/unsolvable labels), train the linear classifier on the training set, and evaluate its classification accuracy on the test set.
The results show that linear classifiers can still separate solvable and unsolvable questions from hidden states, although with reduced accuracy compared to aggregated benchmarks setting (Table 2). This drop is small on HLE but larger on GSM8K. The reason is that HLE contains a larger proportion of unsolvable questions, enabling the linear classifiers to better learn the hidden state information associated with unsolvability. Consequently, these unsolvable questions in HLE can be effectively leveraged to develop models’ Hard Abstention (HA) capability.
Cross-benchmark test. In the above analysis, we think the unsolvable questions in HLE can be effectively leveraged to develop models’ Hard Abstention (HA) capability. Therefore, we train linear classifiers on HLE and evaluate models’ Hard Abstention capability on GSM8k and AIME to test whether this capability generalizes across benchmarks. As shown in Table 7, linear classifiers trained on HLE are able to identify nearly all unsolvable questions in GSM8K and AIME, achieving HA close to 100%.
| Model | GSM8k | AIME |
| GPT-oss-20B | 0 | 0 |
| +Monitor | 100 | 100 |
| R1-Distill-Qwen3-8B | 0 | 0 |
| +Monitor | 100 | 100 |
| R1-Distill-Qwen3-32B | 0 | 0 |
| +Monitor | 90.5 | 94.7 |
| QwQ-32B | 0 | 0 |
| +Monitor | 98.9 | 100 |
In contrast, training on GSM8K is insufficient for developing Hard Abstention capability, because GSM8K contains very few unsolvable questions. The linear classifiers cannot reliably learn the hidden state patterns associated with unsolvability.
Appendix F Ablation Study on Output Prefix
During our initial exploration, we experiment with several output prefixes and select the one that achieves the largest reduction in token usage. We conduct an ablation study comparing:
Initial prefix (our early attempt): Our goal is to prompt the model to acknowledge that it cannot solve an unsolvable question and then provide only a concise potential approach. However, we unexpectedly found that this cannot minimize token usage. The model’s instruction-following ability was insufficient for reliably producing only a concise approach, it often generated overly long responses. The initial prefix is shown below:
Emphasis prefix (final choice): Inspired by prior work (Zhang et al., 2024; Zhao et al., 2021) which show that the final sentence in the instruction strongly influences model behavior, we appended Step 1: at the end of the instruction to explicitly trigger step-by-step approach. Also, we emphasize in the end of the <think>...</think> part that I must give a concise outline to the user (less than 10 steps)!. This combination elicits short, structured responses and lead to the largest reduction in token usage. This is the output prefix we ultimately choose:
Token-budget prefix (following prior work): Following prior work (Han et al., 2025) which compresses unnecessary reasoning by including a reasonable token budget in the prompt, we replace less than 10 steps with an explicit budget constraint use less than 500 tokens. The token-budget prefix is:
| Output prefix | Token |
| No output prefix | 5096 |
| Initial | 4423 |
| Emphasis | 2227 |
| Token-budget | 2567 |
To avoid the confounding effect of incomplete answers, we set the context length to the maximum according to Huggingface model card.
Table 8 shows that all three designs reduce token usage, with the Emphasis prefix achieving the largest reduction. While it is possible that an even better prompt design exists, this does not affect our core conclusion: setting an output prefix enables the model to provide more efficient and more reliable responses on unsolvable questions.
Appendix G Analysis on Coding Task
| Model | Reasoning Expressions | Hidden States | ||
| ConfDiff | ConfCurv | LDA | LR | |
| GPT-oss-20B | 100 | 86.8 | 81.1 | 83.5 |
| R1-Distill-Qwen3-8B | 97.3 | 93.0 | 92.0 | 93.9 |
| R1-Distill-Qwen3-32B | 100 | 91.7 | 81.6 | 84.9 |
| QwQ-32B | 100 | 92.5 | 77.3 | 85.4 |
Since we capture unsolvability signals from reasoning expressions (Section 4) and hidden states (Section 5), without relying on task-specific heuristics, they might in principle generalize to other forms of reasoning tasks. We provide an analysis on a coding task in this section.
We conduct experiments on 1055 samples from LiveCodeBench (Jain et al., 2024). Table 9 shows the accuracy of separating solvable and unsolvable coding questions based on capability boundaries in reasoning expressions and hidden states.
We observe that coding tasks also exhibit clear capability boundaries in both reasoning expressions and hidden states. However, the accuracy from hidden states is lower than that from mathematical tasks. A plausible explanation is that the difficulty of coding problems is less explicitly signaled in the input itself. Unlike in math exams where humans can often immediately recognize “end-of-exam” questions that they are unlikely to solve, coding tasks usually require some degree of initial reasoning before the model can detect unsolvability. As a result, the unsolvability signal in the hidden states of the last input token is weaker, whereas the signal becomes more apparent in the early reasoning expressions.
Appendix H Capability Boundary in Hidden States
@@
Figure 8 visualizes the classification of solvable and unsolvable questions of all LRMs, along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train–test splits. These results indicate that the hidden states of the last input token contain useful signals related to the operational capability boundary, and is valid across different LRMs.
Appendix I Monitoring Strategies with Maximum Context Length
To evaluate whether our monitoring strategies (Monitor and Monitor) hold promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024), we provide the optimized token usage under extended token budget, set to the maximum according to HuggingFace model card (128k for GPT-oss-20B, 64k for R1-Distill-Qwen3-8B, and 32k for both R1-Distill-Qwen3-32B and QwQ-32B).
| Context Length | Token ↓ | Overflow (%) ↓ | |
| GPT-oss-20B | 128k | ||
| +Monitor | |||
| +Monitor | |||
| R1-Distill-Qwen3-8B | 64k | ||
| +Monitor | |||
| +Monitor | |||
| R1-Distill-Qwen-32B | 32k | ||
| +Monitor | |||
| +Monitor | |||
| QwQ-32B | 32k | ||
| +Monitor | |||
| +Monitor |
Under these settings, Table 10 shows that the models always produce complete answers (Overflow = 0), and our monitoring strategies still reduce token usage significantly as expected (Acc and HA are the same as under 2k or 4k context length).
Moreover, it is worth noting that our strategies are not constrained by context length:
Monitor continuously tracks the densities of confident and uncertain expressions throughout the reasoning process and can detect unsolvable questions at an early stage (Figure 3), well before the model reaches the end of reasoning.
Monitor identifies unsolvable questions directly from the hidden states of last input token, i.e., even before the model begins reasoning.
Because our strategies operate early in the reasoning trajectory, we intentionally used relatively short context length (e.g., 2k or 4k) in Section 6 to demonstrate their early detection capabilities.
Appendix J LLMs Usage
In this work, LLMs are used to polish part of the writing and to revise portions of the code. All core contributions are carried out by the authors, including the main ideas, literature review, experimental design and execution, figure and table preparation, and analysis.
Appendix K Unproductive Reasoning
@@ Appendix I Monitoring Strategies with Maximum Context Length
Figure 9 and Figure 10 show representative unproductive reasoning phenomena in LRMs, including repetitive looping and error accumulation, which impair their capacity for sound and consistent reasoning in complex problem-solving scenarios. They respectively show how models get stuck in repetitive deductions (Figure 9) and how errors compound during formula application and result derivation (Figure 10). In mathematical problem reasoning, the former presents repeated deduction of the same uncertain conclusion, while the latter sees continuous error accumulation due to flawed initial methods.